OpenStudy (anonymous):

For the function defined by: f(x)={x^2, x<= 1 {2x+1, x>1 Evaluate f(0)

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

I'm not sure I tried f(0)={0^2, 0<=1 {0+1, 0>1 and then I didn't know what to do from there

OpenStudy (anonymous):

it is not clear to you how to read the function

OpenStudy (anonymous):

it means \(f(x)=x^2\) IF \(x\leq 1\) and \(f(x)=2x+1\) IF \(x>1\)

OpenStudy (anonymous):

No it isn't clear

OpenStudy (anonymous):

so if i want for example \(f(5)\) first i ask whether 5 is larger than 1 or smaller than 1

OpenStudy (anonymous):

since \(5>1\) i use the second formula and \(f(5)=2\times 5+1=11\)

OpenStudy (anonymous):

whereas if i want \(f(-2)\) i see that \(-2\) is less than \(1\) so i use the first formula and \(f(-2)=(-2)^2=4\)

OpenStudy (anonymous):

Ohh Okay, I'm understanding a little bit more now

OpenStudy (anonymous):

\[ f(x) = \left\{ \begin{array}{lr} x^2 & \text{if } x\leq 1\\ 2x+1 &\text{ if }x>1 \end{array} \right.\]

OpenStudy (anonymous):

missing the word "if" so if you want \(f(0)\) since \(0\) is less than 1, use the top formula

OpenStudy (anonymous):

so what do I do with the bottom formula ?

OpenStudy (anonymous):

How would I write the final answer? @satellite73

OpenStudy (anonymous):

would it be f(0)=0< 1